OpenStudy (anonymous):

The label on a 2-liter bottle of soda indicates that one serving is 8 fluid ounces and that there are 8 servings in the bottle. Are these exact measurements? Explain, using mathematics to justify your answer.

OpenStudy (tootsi123):

what kind of math are you working on like geometry, trig, etc.

OpenStudy (anonymous):

YEA geometry

OpenStudy (tootsi123):

ugh i hate geometry i'm doing it to lol are u online with it or nah

OpenStudy (anonymous):

yea

OpenStudy (tootsi123):

08 servings of 8 fl oz = 64 fl oz 1 gal = 3.785411784 liters (exact) 1 gal = 128 fl oz (exact) 64 fl oz * (1 gal/128 fl oz) * (3.785 L/1 gal) = 1.89 L 8 servings of 8 fl oz are approximately 1.89 liters, not 2 liters 51 03.62 ounces** left over me to XD on odessy i hate it ugh but yeah theres your answer

OpenStudy (anonymous):

how about this? One serving of water is considered 8 fluid ounces. It is recommended that adults drink 8 servings of water per day. How many quarts of water is this? 1 quart 2 quarts 4 quarts 8 quarts

OpenStudy (tootsi123):

alright

OpenStudy (tootsi123):

2 quarts XD

OpenStudy (anonymous):

1 quart is 32 fl oz so if 8 oz is a serving and there are 8 servings 8*8 is 64 oz. 64 oz/32 oz= 2 quarts…

OpenStudy (anonymous):

Btw Hippo :D I'm in Geometry too :D

OpenStudy (anonymous):

The average hair growth rate is one-half inch per month. Express this rate in millimeters per month. 1 point two seven millimeters per month 3 point eight one millimeters per month twelve point seven millimeters per month thirty-eight point one millimeters per month

OpenStudy (tootsi123):

1 inch = 25.4 mm one-half inch = 25.4/2